# Generalizing Reward Modeling for Out-of-Distribution Preference Learning

### v1

@@ 4.2 Evaluation with Preference Data
(a) Training PL loss and accuracy on held-out domain: book.
(b) Evaluation PL accuracy on held-out domain: book.
(c) Training PL loss and accuracy on held-out domains: ac&an&ba.
(d) Evaluation PL accuracy on held-out domains: ac&an&ba.
Figure 2: PL accuracy $\mathcal{A}_{\rm PL}$ against the training steps.
Preference data from the test distributions is utilized for evaluation. Specifically, we examine the optimization of the outer objective against the training steps, as depicted in Figure 2. Figure 2 (a) and (c) demonstrate that the PL loss $\ell_{\rm PL}$ decreases rapidly within the initial 600 steps for the controlled sentiment generation task and within the initial 3,000 steps for the knowledge answer generation task, respectively. Subsequently, the PL loss reaches a low value for both tasks, exhibiting only minor fluctuations. The PL accuracy $\mathcal{A}_{\rm PL}$ increases in the opposite trend to the PL loss and eventually surpasses $0.9$ for both tasks. This indicates that our method is effective in conducting in-distribution preference learning. To compare various methods for out-of-distribution preference learning, Figure 2 (b) and (d) depict the PL accuracy $\mathcal{A}_{\rm PL}$ on the evaluation set of test distributions for controlled sentiment generation and knowledge answer generation, respectively. In comparison to the SFT baseline, other methods utilize preference data for training, resulting in higher PL accuracy for both tasks. Moreover, our method optimizes a bilevel optimization objective for out-of-distribution preference learning, significantly outperforming other methods by achieving a PL accuracy of $0.86$ for controlled sentiment generation and $0.75$ for knowledge answer generation.
@@ 4.3 Evaluation with a Learned RM
(a) Held-out domain: book.
(b) Held-out domain: dvd.
(c) Held-out domain: electronics.
(d) Held-out domain: kitchen.
Figure 3: Reward on the amazon review sentiment generation.
[…]
@@ 4.4 Evaluation with GPT-4 Judgement
target domain ac,an,ba ca,cu,do ne,hi,hr
ac an ba avg. ca cu do avg. ne hi hr avg.
Base Model 0.153 0.234 0.198 0.195 0.225 0.124 0.181 0.177 0.176 0.206 0.134 0.132
Sft 0.587 0.543 0.495 0.542 0.478 0.521 0.576 0.525 0.578 0.525 0.456 0.520
Ppo (tr) 0.646 0.586 0.557 0.596 0.507 0.604 0.558 0.556 0.660 0.557 0.502 0.573
Ppo (tr+te) 0.682 0.594 0.567 0.614 0.557 0.614 0.583 0.585 0.672 0.587 0.536 0.598
Dpo 0.657 0.554 0.565 0.592 0.525 0.537 0.612 0.558 0.648 0.506 0.527 0.560
Hpl 0.536 0.497 0.624 0.552 0.556 0.625 0.542 0.574 0.682 0.486 0.478 0.549
our method 0.724 0.607 0.724 0.685 0.686 0.687 0.702 0.692 0.686 0.652 0.586 0.642
target domain phi,phy,sci scif,so,ve ch,ex,le
phi phy sci avg. scif so ve avg. ch ex le avg.
Base Model 0.152 0.237 0.167 0.185 0.188 0.193 0.215 0.199 0.203 0.113 0.124 0.147
Sft 0.467 0.506 0.456 0.476 0.526 0.497 0.502 0.508 0.546 0.487 0.568 0.534
Ppo (tr) 0.568 0.545 0.563 0.558 0.602 0.507 0.478 0.529 0.654 0.604 0.625 0.628
Ppo (tr+te) 0.601 0.552 0.597 0.583 0.640 0.528 0.601 0.590 0.647 0.654 0.639 0.647
Dpo 0.507 0.582 0.557 0.549 0.580 0.527 0.554 0.554 0.636 0.452 0.657 0.582
Hpl 0.342 0.478 0.572 0.467 0.582 0.486 0.557 0.542 0.486 0.627 0.586 0.566
our method 0.604 0.686 0.724 0.671 0.656 0.567 0.652 0.625 0.672 0.680 0.749 0.700
Table 3: GPT-4 evaluation of win rate on the knowledge answer generation. We format the best, the second best and worse than SFT results.
Building on previous work, we utilize GPT-4 to evaluate whether an answer generated by the model is more correct and useful compared to the preferred response from the knowledge answer generation dataset. We then assess the win rates of all methods against the preferred answers in the dataset. As depicted in Table 3, all fine-tuning methods significantly outperform the base model of Flan-T5 XXL. This highlights the importance of fine-tuning a pretrained language model with supervised data for the knowledge answer generation task. In comparison to the Sft baseline, the other methods utilize human preference data for further fine-tuning, resulting in better win rates across all held-out domains. Compared to Ppo (tr), Ppo (tr+te) utilizes supervised data from test distributions for fine-tuning and achieves higher average win rates for each experiment, demonstrating the effectiveness of using in-distribution data for fine-tuning a language model. However, all baselines including Ppo, Dpo, and Hpl are constrained by direct training on in-distribution preference data, and thus can only learn in-distribution preference information. In contrast, we employ a bilevel optimization approach to learn a general reward function, which can be applied to out-of-distribution policy fine-tuning. Consequently, our method achieves the best results across all out-of-distribution experiments.
@@ 4.5 Effects of $\beta$ on Reward Learning
(a) Reward on meta-test set of book.
(b) logarithm of combination term on meta-test set of book.
(c) PL accuracy on test set of book.
(d) Reward on meta-test set of ac,an,ba.
(e) logarithm of combination term on meta-test set of ac,an,ba.
(f) PL accuracy on test set of ac,an,ba.
Figure 4: The effects of reward learning w.r.t. the reward controlling factor $\beta$.
Figure 4 illustrates our investigation into various values of $\beta\in{0.1,1.0,2.0,4.0,8.0}$ for reward learning. From Figures 4 (a) and (d), it is evident that when the reward controlling factor is too small, such as $\beta=0.1$, the reward distribution becomes very sharp and approaches 00, providing limited information for policy fine-tuning. As $\beta$ increases, the reward distribution becomes flatter, which better represents the preferences of inputs for guiding policy fine-tuning. Additionally, we explore the learning of the combination term $\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)$ for cross-entropy in the inner objective. Figures 4 (b) and (e) demonstrate that as $\beta$ increases, the distribution of the combination term approaches a standard normal distribution, facilitating faster convergence of policy fine-tuning. Furthermore, we examine how $\beta$ influences preference learning by presenting the preference learning accuracy against the training steps for various values of $\beta$ in Figures 4 (c) and (f). We observe that for each value of $\beta$, as the number of training steps increases, the preference learning accuracy improves and eventually converges to a high plateau. Moreover, a larger value of $\beta$ yields better preference learning accuracy for both tasks. The experimental findings align with the theoretical result of Theorem 1, suggesting that a larger number of iterations $K$ and a larger value of $\beta$ contribute to the algorithm’s convergence to a more accurate stationary point.
@@ 5 Conclusion
Preference learning (PL) is a powerful framework for training language models to align with human values. Instead of discretely training a reward model solely for in-distribution PL, this work investigated an out-of-distribution (OOD) PL algorithm. The aim is to improve the generalization of reward modeling for aligning a language policy with human preferences across various test distributions. To address OOD PL, we proposed a meta-learning approach conducted through a bilevel optimization algorithm. The outer objective aims to align the predicted probability generated by the language policy with human preferences, while the inner objective focuses on performing regularized policy optimization with the learned reward model. As a result, when encountering a novel test distribution, fine-tuning the language policy using only supervised data can achieve alignment with human preferences. By deriving the inner objective, we revised the regularization term in previous work on RLHF by incorporating consideration of the target distribution to mitigate policy drift. Additionally, we employed a parameter estimation approach to reduce the RL objective to a regression problem for more efficient training. We derived an upper bound for the convergence rate of reward model learning using the proposed bilevel optimization algorithm. This bound demonstrates significant relationships with the number of outer-loop iterations and the controlling factor of reward modeling. We conducted experiments on both controlled sentiment generalization and knowledge answer generation. Evaluation with preference data demonstrates that the generalized reward modeling achieved by our meta-learning approach effectively guides policy learning to align with human preferences. Evaluation on sentiment generation with a learned reward model indicates that our method achieves superior rewards compared to other OOD PL baselines across four domains. Evaluation using GPT-4 judgment demonstrates that our method outperforms a range of strong baselines, achieving the highest win rate across 16 held-out domains.
Appendix A Parameter Estimation for Policy Learning
@@
$\mathop{\arg\max}_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi(\cdot|x)}\left[r_{\phi}(x,y)\right]-\beta\mathrm{D}_{\rm KL}\left(\pi(\cdot|x)||\hat{\mu}(\cdot|x)\right)\right]$
$=\mathop{\arg\max}_{\pi}\int_{x}\mathcal{D}(x)\left(\int_{y}\pi(y|x)r_{\phi}(x,y)dy-\beta\mathrm{D}_{\rm KL}\left(\pi(\cdot|x)||\hat{\mu}(\cdot|x)\right)\right)dx$
@@
$\mathop{\arg\max}_{\pi}$ $\int_{x}\mathcal{D}(x)\left(\int_{y}\pi(y|x)r_{\phi}(x,y)dy-\beta\mathrm{D}_{\rm KL}\left(\pi(\cdot|x)||\hat{\mu}(\cdot|x)\right)\right)dx$
@@
$\mathcal{L}(\pi,\alpha)=\int_{x}\mathcal{D}(x)\left(\int_{y}\pi(y|x)r_{\phi}(x,y)dy-\beta\mathrm{D}_{\rm KL}\left(\pi(\cdot|x)||\hat{\mu}(\cdot|x)\right)\right)dx+\int_{x}\alpha_{x}\left(1-\int_{y}\pi(y|x)dy\right)dx,$
@@
$\pi_{*}(y|x)=\hat{\mu}(y|x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)\exp\left(-\frac{\alpha_{x}}{\beta\mathcal{D}(x)}-1\right),\quad\forall y\forall x$ (27)
@@
$\frac{\partial^{2}\mathcal{L}(\pi,\alpha)}{\partial[\pi(y|x)]^{2}}=\int_{x}\int_{y}-\frac{\beta\mathcal{D}(x)}{\pi(y|x)}dydx<0,$
since $\beta,\mathcal{D}(x),\pi(y|x)>0$. This indicates the strict concaveness of $\mathcal{L}(\pi,\alpha)$. Therefore, Eq. (27) indicates a global optimization solution $\pi_{*}=\mathop{\arg\max}_{\pi}\mathcal{L}(\pi,\alpha)$.
@@
$\pi^{*}(y|x)=\frac{1}{Z(x)}\hat{\mu}(y|x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right),\quad\forall y\forall x$ (28)
To obtain such a global optimization solution for the policy, we use a parameter estimation method, i.e., using a machine learning method to train a parameter estimated policy $\pi_{\theta}$ close to the globally optimal solution $\pi^{*}$ of Eq. (28) by solving the following supervised regression problem:
$\mathop{\arg\min}_{{\theta}}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathrm{D}_{\rm KL}\left(\pi^{*}(\cdot|x)||\pi_{\theta}(\cdot|x)\right)\right]$
$=$ $\mathop{\arg\min}_{{\theta}}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathrm{D}_{\rm KL}\left(\frac{1}{Z(x)}\hat{\mu}(y|x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)\bigg{\|}\pi_{\theta}(\cdot|x)\right)\right]$
$=$ $\mathop{\arg\min}_{\theta}-\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\hat{\mu}(y|x)}\left[\log\pi_{\theta}(y|x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)\right]$
Appendix B Proof of Proposition 1
@@
$\frac{\partial\ell_{\rm PL}({\phi_{k}},{\theta_{k,D}})}{\partial{\phi}_{k}}$ $=-\frac{\partial}{\partial{\phi}_{k}}\log\sigma\left(\log\pi_{\theta_{k,D}}(y_{k}|x_{k})-\log\pi_{\theta_{k,D}}(y^{\prime}_{k}|x_{k})\right)$ (29)
$=-\frac{\partial{\theta}_{k,D}}{\partial{\phi}_{k}}\left(\nabla_{{\theta}}\log\pi_{\theta_{k,D}}(y_{k}|x_{k})-\nabla_{{\theta}}\log\pi_{{\theta}_{k,D}}(y^{\prime}_{k}|x_{k})\right)$ (30)
@@
$=$ ${\theta}_{k,t}+\alpha\nabla_{\theta}\left(\log\pi_{\theta_{k,t}}(y_{k,t}|x_{k,t})\exp\left(\frac{1}{\beta}r_{\phi_{k}}(x_{k,t},y_{k,t})\right)\right)$
@@
$+\frac{\alpha}{\beta}\nabla_{{\theta}}\log\pi_{\theta_{k,t}}(y_{k,t}|x_{k,t})\exp\left(\frac{1}{\beta}r_{\phi_{k}}(x_{k,t},y_{k,t})\right)$
To simplify the presentation, let $R^{k,t}:=\exp\left(\frac{1}{\beta}r_{\phi_{k}}(x_{k,t},y_{k,t})\right)$, $F^{k,t}:=\log\pi_{\theta_{k,t}}(y_{k,t}|x_{k,t})$ and $\nabla:=\nabla_{{\theta}}$, we have
$\frac{\partial{\theta}_{k,t+1}}{\partial{\phi}_{k}}=$ $\frac{\partial{\theta}_{k,t}}{\partial{\phi}_{k}}\left(I+\alpha R^{k,t}\nabla^{2}F^{k,t}\right)+\frac{\alpha}{\beta}R^{k,t}\nabla F^{k,t}$
Telescoping the above equality over $t$ from $1$ to $D$ yields
$\begin{split}\frac{\partial{\theta}_{k,D}}{\partial{\phi_{k}}}=&\frac{\partial{\theta}_{k,0}}{\partial{\phi_{k}}}\prod_{t=0}^{D-1}\left(I+\alpha R^{k,t}\nabla^{2}F^{k,t}\right)+\frac{\alpha}{\beta}\sum_{t=0}^{D-1}R^{k,t}\nabla F^{k,t}\prod_{j=t+1}^{D-1}\left(I+\alpha R^{k,t}\nabla^{2}F^{k,t}\right)\\ =&\frac{\alpha}{\beta}\sum_{t=0}^{D-1}R^{k,t}\nabla F^{k,t}\prod_{j=t+1}^{D-1}\left(I+\alpha R^{k,t}\nabla^{2}F^{k,t}\right),\end{split}$ (31)
where the second equation follows from the fact that $\frac{\partial{\theta}_{k,0}}{\partial\phi_{k}}=0$. Combining Eq. (29) and Eq. (31) finishes the proof.
Appendix C Supporting Lemmas
Lemma 1.
Based on Assumption 1, 2 and 3, $\mathcal{L}_{\rm FT}(\phi,\theta)$ is $\frac{BL}{\exp\beta}$-smoothness and $\frac{b\mu}{\exp\beta}$-strong convexity w.r.t. $\theta$.
Proof.

### v2

@@ 3.6 Evaluation with Preference Data
(a) book.
(b) dvd.
(c) electronics.
(d) kitchen.
(e) book.
(f) dvd.
(g) electronics.
(h) kitchen.
Figure 3: (a)-(d) illustrate the training loss and PL accuracy $\mathcal{A}_{\rm PL}$ domain, and (e)-(h) illustrate the evaluation accuracy against the training steps for each held-out CSG distribution.
Preference data from the test distributions is utilized for evaluation. Specifically, we examine the optimization of the outer objective against the training steps, as depicted in Fig. 3. (a) and (c) demonstrate that the PL loss $\ell_{\rm PL}$ decreases rapidly within the initial 600 steps for the CSG task and within the initial 3,000 steps for the KAG task, respectively. Subsequently, the PL loss reaches a low value for both tasks, exhibiting only minor fluctuations. The PL accuracy $\mathcal{A}_{\rm PL}$ increases in the opposite trend to the PL loss and eventually surpasses $0.9$ for both tasks. This indicates that our method is effective in conducting ID PL. To compare various methods for OOD PL, Fig. 3 (b) and (d) depict the PL accuracy $\mathcal{A}_{\rm PL}$ on the evaluation set of test distributions for CSG and KAG, respectively. In comparison to the SFT baseline, other methods utilize preference data for training, resulting in higher PL accuracy for both tasks. Moreover, our method optimizes a bilevel optimization objective for OOD PL, significantly outperforming other methods by achieving a PL accuracy of $0.86$ for CSG and $0.75$ for KAG.
@@ 3.7 Evaluation with a Learned RM
(a) book.
(b) dvd.
(c) electronics.
(d) kitchen.
Figure 4: Reward on four held-out Amazon Review distributions.
Following the approach of Rafailov et al. [30], we employ a learned RM to evaluate PL in CSG. Specifically, we fine-tune a RoBERTaLARGE model on four domains of the Amazon Review dataset to serve as the reference reward function. As depicted in Fig. 4, we compare various OOD PL methods across four held-out domains, using different temperatures ranging from ${0.25,0.5,0.75,1.0}$ to control the randomness of generation. The results for each temperature represent the average of three repetitions of generation. From Fig. 4, it is observed that as the temperature increases from $0.25$ to $1.0$, the variances of rewards for all methods tend to increase across the four held-out domains. Despite the wide range of rewards observed for all methods, our method consistently outperforms other baselines in terms of median reward across all temperatures. This demonstrates the effectiveness of our proposed approach in optimizing a general RM for OOD PL.
@@ 3.8 Evaluation with GPT-4 Judgement
Table 3: GPT-4 evaluation of win rate on the KAG. We format the best, the second best and worse than SFT results.
Method ac,an,ba ca,cu,do ne,hi,hr phi,phy,sci scif,so,ve ch,ex,le
ac an ba avg. ca cu do avg. ne hi hr avg. phi phy sci avg. scif so ve avg. ch ex le avg.
Base Model .153 .234 .198 .195 .225 .124 .181 .177 .176 .206 .134 .132 .152 .237 .167 .185 .188 .193 .215 .199 .203 .113 .124 .147
SFT .627 .583 .535 .582 .498 .541 .596 .545 .598 .545 .476 .540 .487 .526 .476 .496 .526 .497 .502 .508 .646 .587 .668 .634
PPO (tr) .646 .586 .557 .596 .507 .604 .558 .556 .660 .557 .502 .573 .538 .515 .533 .528 .602 .507 .478 .529 .654 .604 .625 .628
PPO (te) .682 .594 .567 .614 .557 .614 .583 .585 .672 .587 .536 .598 .601 .552 .597 .583 .640 .528 .601 .590 .647 .654 .639 .647
DPO .657 .554 .565 .592 .525 .537 .612 .558 .648 .506 .527 .560 .507 .582 .557 .549 .580 .527 .554 .554 .636 .452 .657 .582
HPL .536 .497 .624 .552 .556 .625 .542 .574 .682 .486 .478 .549 .342 .478 .572 .467 .582 .486 .557 .542 .486 .627 .586 .566
Ours .724 .607 .724 .685 .686 .687 .702 .692 .686 .652 .586 .642 .604 .686 .724 .671 .656 .567 .652 .625 .672 .680 .749 .700
Building upon prior research [30], we employ GPT-4 to assess whether the model-generated answers are better than the preferred responses from the SHP dataset. Subsequently, we evaluate the win rates of all approaches against the preferred answers in the test set. This underscores the significance of fine-tuning a pretrained language model using supervised data for KAG. Compared to the SFT baseline, PPO (tr) additionally utilizes the training distributions for RL fine-tuning but struggles to enhance the results on the test distribution. This illustrates the challenge of OOD PL with RLHF. Additionally, DPO and HPL fine-tune the SFT model with PL data from the training distributions but also struggle to enhance the OOD results. The limitations are evident in the distribution bias of RM training. In contrast, we employ a meta-learning approach to learn a general RM that can be utilized for OOD policy tuning. Consequently, our method achieves superior results across all experiments.
@@ 3.9 Effects of $\beta$ on Reward Learning
(a) Reward.
(b) log. com.
(c) PL acc.
(d) Reward.
(e) log. com.
(f) PL acc.
Figure 5: Effects of reward learning w.r.t. the reward controlling factor $\beta$ on the book distribution (a-c) and on the ac,an,ba distribution (d-f).
Fig. 5 illustrates our investigation into various values of $\beta\in{0.1,1.0,2.0,4.0,8.0}$ for reward learning. From Fig. 5 (a) and (d), it is evident that when the reward controlling factor is too small, such as $\beta=0.1$, the reward distribution becomes very sharp and approaches 00, providing limited information for policy fine-tuning. As $\beta$ increases, the reward distribution becomes flatter, which better represents the preferences of inputs for guiding policy fine-tuning. Additionally, we explore the learning of the combination term $\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)$ for cross-entropy in the inner objective. Figures 5 (b) and (e) demonstrate that as $\beta$ increases, the distribution of the combination term approaches a standard normal distribution, facilitating faster convergence of policy fine-tuning. Furthermore, we examine how $\beta$ influences PL by presenting the PL accuracy against the training steps for various values of $\beta$ in Fig. 5 (c) and (f). We observe that for each value of $\beta$, as the number of training steps increases, the PL accuracy improves and eventually converges to a high plateau. Moreover, a larger value of $\beta$ yields better PL accuracy for both tasks. The experimental findings align with the theoretical result of Theorem 3.1, suggesting that a larger number of iterations $K$ and a larger value of $\beta$ contribute to the algorithm’s convergence to a more accurate stationary point.
@@ 4 Conclusion
Preference learning (PL) is a powerful framework for training language models to align with human values. Instead of discretely training a reward model solely for in-distribution (ID) PL, this work investigated an out-of-distribution (OOD) PL algorithm. The aim is to improve the generalization of reward modeling across various distributions. To address OOD PL, we proposed a meta-learning approach conducted through a bilevel optimization algorithm. We derived an upper bound for the convergence rate of the proposed algorithm. We conducted experiments on both controlled sentiment generation and knowledge answer generation. Results demonstrate that the generalized reward modeling achieved by our meta-learning approach effectively guides policy learning to align with human preferences.
Appendix 0.A Parameter Estimation for Policy Learning
@@
$\mathop{\arg\max}_{\pi}\mathbb{E}_{x\sim\mathcal{D}_{x}}\left[\mathbb{E}_{y\sim\pi(\cdot|x)}\left[r_{\phi}(x,y)\right]-\beta\mathrm{D}_{\rm KL}\left(\pi(\cdot|x)||\mathcal{D}_{y|x}(\cdot)\right)\right]$
$=\mathop{\arg\max}_{\pi}\int_{x}\mathcal{D}_{x}(x)\left(\int_{y}\pi(y|x)r_{\phi}(x,y)dy-\beta\mathrm{D}_{\rm KL}\left(\pi(\cdot|x)||\mathcal{D}_{y|x}(\cdot)\right)\right)dx$
@@
$\mathop{\arg\max}_{\pi}$ $\int_{x}\mathcal{D}_{x}(x)\left(\int_{y}\pi(y|x)r_{\phi}(x,y)dy-\beta\mathrm{D}_{\rm KL}\left(\pi(\cdot|x)||\mathcal{D}_{y|x}(\cdot)\right)\right)dx$
@@
$\mathcal{L}(\pi,\alpha)=$ $\int_{x}\mathcal{D}_{x}(x)\left(\int_{y}\pi(y|x)r_{\phi}(x,y)dy-\beta\mathrm{D}_{\rm KL}\left(\pi(\cdot|x)||\mathcal{D}_{y|x}(\cdot)\right)\right)dx$
[…]
@@
$\pi_{*}(y|x)=\mathcal{D}_{y|x}(y)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)\exp\left(-\frac{\alpha_{x}}{\beta\mathcal{D}(x)}-1\right),\quad\forall y\forall x$ (24)
@@
$\frac{\partial^{2}\mathcal{L}(\pi,\alpha)}{\partial[\pi(y|x)]^{2}}=\int_{x}\int_{y}-\frac{\beta\mathcal{D}_{x}(x)}{\pi(y|x)}dydx<0,$
since $\beta,\mathcal{D}_{x}(x),\pi(y|x)>0$. This indicates the strict concaveness of $\mathcal{L}(\pi,\alpha)$. Therefore, Eq. (24) indicates a global optimization solution $\pi_{*}=\mathop{\arg\max}_{\pi}\mathcal{L}(\pi,\alpha)$.
@@
$\pi_{*}(y|x)=\frac{1}{Z(x)}\mathcal{D}_{y|x}(y)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right),\quad\forall y\forall x$ (25)
To obtain such a global optimization solution for the policy, we use a parameter estimation method, i.e., using a machine learning method to train a parameter estimated policy $\pi_{\theta}$ close to the globally optimal solution $\pi_{*}$ of Eq. (25) by solving the following supervised regression problem:
$\mathop{\arg\min}_{{\theta}}\mathbb{E}_{x\sim\mathcal{D}_{x}}\left[\mathrm{D}_{\rm KL}\left(\pi_{*}(\cdot|x)||\pi_{\theta}(\cdot|x)\right)\right]$
$=$ $\mathop{\arg\min}_{{\theta}}\mathbb{E}_{x\sim\mathcal{D}_{x}}\left[\mathrm{D}_{\rm KL}\left(\frac{1}{Z(x)}\mathcal{D}_{y|x}(\cdot)\exp\left(\frac{1}{\beta}r_{\phi}(x,\cdot)\right)\bigg{\|}\pi_{\theta}(\cdot|x)\right)\right]$
$=$ $\mathop{\arg\min}_{\theta}-\mathbb{E}_{x\sim\mathcal{D}_{x}}\mathbb{E}_{y\sim\mathcal{D}_{y|x}}\left[\log\pi_{\theta}(y|x)\exp\left(\frac{1}{\beta}r_{\phi}(x,y)\right)\right]$
Appendix 0.B Proof of Proposition 1
@@
$\frac{\partial\ell_{\rm PL}({\phi_{k}},{\theta_{k,D}})}{\partial{\phi}_{k}}$
$=$ $-\frac{\partial}{\partial{\phi}_{k}}\log\sigma\left(\log\pi_{\theta_{k,D}}(y_{k}|x_{k})-\log\pi_{\theta_{k,D}}(y^{\prime}_{k}|x_{k})\right)$
$=$ $-\frac{\partial{\theta}_{k,D}}{\partial{\phi}_{k}}\left(\nabla_{{\theta}}\log\pi_{\theta_{k,D}}(y_{k}|x_{k})-\nabla_{{\theta}}\log\pi_{{\theta}_{k,D}}(y^{\prime}_{k}|x_{k})\right)$
$\cdot\left(1-\sigma\left(\log\pi_{\theta_{k,D}}(y_{k}|x_{k})-\log\pi_{\theta_{k,D}}(y^{\prime}_{k}|x_{k})\right)\right)$
Then, let $A_{k}=\log\pi_{\theta_{k,D}}(y_{k}|x_{k})-\log\pi_{{\theta}_{k,D}}(y^{\prime}_{k}|x_{k})$, we have
$\frac{\partial\ell_{\rm PL}({\phi_{k}},{\theta_{k,D}})}{\partial{\phi}_{k}}=-\frac{\partial{\theta}_{k,D}}{\partial{\phi}_{k}}\nabla_{{\theta}}A_{k}\left(1-\sigma\left(A_{k}\right)\right)$ (26)
@@
$+\frac{\alpha}{\beta}\nabla_{\phi}r_{\phi_{k}}(x_{k,t},y_{k,t})\nabla_{{\theta}}^{\top}\log\pi_{\theta_{k,t}}(y_{k,t}|x_{k,t})\exp\left(\frac{1}{\beta}r_{\phi_{k}}(x_{k,t},y_{k,t})\right)$
To simplify the presentation, let $R_{k,t}:=\exp\left(\frac{1}{\beta}r_{\phi_{k}}(x_{k,t},y_{k,t})\right)$, $F_{k,t}:=\log\pi_{\theta_{k,t}}(y_{k,t}|x_{k,t})$ and $H_{k,t}:=r_{\phi_{k}}(x_{k,t},y_{k,t})$, we have
$\frac{\partial{\theta}_{k,t+1}}{\partial{\phi}_{k}}=$ $\frac{\partial{\theta}_{k,t}}{\partial{\phi}_{k}}\left(I+\alpha\nabla_{{\theta}}^{2}F_{k,t}R_{k,t}\right)+\frac{\alpha}{\beta}\nabla_{\phi}H_{k,t}\nabla_{\theta}^{\top}F_{k,t}R_{k,t}$
Telescoping the above equality over $t$ from 00 to $D-1$ yields
$\begin{split}\frac{\partial{\theta}_{k,D}}{\partial{\phi_{k}}}=&\frac{\partial{\theta}_{k,0}}{\partial{\phi_{k}}}\prod_{t=0}^{D-1}\left(I+\alpha\nabla_{{\theta}}^{2}F_{k,t}R_{k,t}\right)+\frac{\alpha}{\beta}\sum_{t=0}^{D-1}\nabla_{\phi}H_{k,t}\nabla_{{\theta}}^{\top}F_{k,t}R_{k,t}\prod_{j=t+1}^{D-1}\left(I+\alpha\nabla_{{\theta}}^{2}F_{k,t}R_{k,t}\right)\\ =&\frac{\alpha}{\beta}\sum_{t=0}^{D-1}\nabla_{\phi}H_{k,t}\nabla_{{\theta}}^{\top}F_{k,t}R_{k,t}\prod_{j=t+1}^{D-1}\left(I+\alpha\nabla_{{\theta}}^{2}F_{k,t}R_{k,t}\right),\end{split}$ (27)
where the second equation follows from the fact that $\frac{\partial{\theta}_{k,0}}{\partial\phi_{k}}=0$. Combining Eq. (0.B) finishes the proof.
Appendix 0.C Supporting Lemmas
Lemma 1
Based on Assumption 1, 2 and 3, $\ell_{\rm FT}(\phi,\theta)$ ($\mathcal{L}_{\rm FT}(\phi,\theta)$) is $\frac{BL}{\exp\beta}$-smoothness and $\frac{b\mu}{\exp\beta}$-strong convexity w.r.t. $\theta$.
Proof